# Assessing AI vs Human-Authored Spear Phishing SMS Attacks: An Empirical Study Using the TRAPD Method

### v1

@@ Abstract
This paper explores the rising concern of utilizing Large Language Models (LLMs) in spear phishing message generation, and their performance compared to human-authored counterparts. Our pilot study compares the effectiveness of smishing (SMS phishing) messages created by GPT-4 and human authors, which have been personalized to willing targets. The targets assessed the messages in a modified ranked-order experiment using a novel methodology we call TRAPD (Threshold Ranking Approach for Personalized Deception). Specifically, targets provide personal information (job title and location, hobby, item purchased online), spear smishing messages are created using this information by humans and GPT-4, targets are invited back to rank-order 12 messages from most to least convincing (and identify which they would click on), and then asked questions about why they ranked messages the way they did. They also guess which messages are created by an LLM and their reasoning. Results from 25 targets show that LLM-generated messages are most often perceived as more convincing than those authored by humans, with messages related to jobs being the most convincing. We characterize different criteria used when assessing the authenticity of messages including word choice, style, and personal relevance. Results also show that targets were unable to identify whether the messages was AI-generated or human-authored and struggled to identify criteria to use in order to make this distinction. This study aims to highlight the urgent need for further research and improved countermeasures against personalized AI-enabled social engineering attacks.
@@ 1 Introduction
In today’s digital landscape, cybersecurity defenders and adversaries continuously adapt to new and emerging technologies, many of which create new cyber risks. While implementing robust security measures can significantly reduce risk posed by potential threat actors, the strength of a system is dependent upon its users [1]. Among the myriad end user vulnerabilities, and tactics employed by malicious actors, phishing remains the most common way of infiltrating systems. Phishing is a social engineering tactic that persuades victims to take an action (e.g., click on a malicious link or email attachment) that causes malicious code to run or discloses sensitive information. Phishing attacks usually take the form of email messages, telephone calls (vishing) or SMS messages (smishing) [2]. Phishing attacks are by far the most numerous, with the FBI Internet Crime Complaint Center reporting five times more phishing complaints in 2023 than any other attack category, totaling over \$300 million in reported complaint losses from 2021 to 2023 [3].
@@
Spear phishing is a targeted form of phishing that involves the use of more personalized messages utilizing specific information about a user (e.g., name, job title, home address) to make the message more believable, thus making it more difficult to differentiate from legitimate sources [4]. Targeted spear phishing poses a significant threat for the future as the method continues to grow. The Proofpoint “State of the Phish 2023” report noted that spear phishing prevalence represented approximately 74% of attacks in 2022, as opposed to bulk phishing at 8%.[5]. In instances where the pretext used against a victim matches their expectation, the attacker is likely to succeed [6]. For example, when a user expects to receive a product purchase and shipping email from a large online retailer and receives a phishing email claiming to provide a shipping update, the user is likely to fall victim [7].
Traditionally, crafting these targeted messages were less common, requiring significant time and effort on the part of the cybercriminals to carry out research and then craft specifically targeted (rather than generic) messages [7]. However, with the advancement of AI, emergent large language models (LLMs) could potentially be used to carry out spear phishing campaigns with greater efficiency and human-like accuracy [8]. This recent shift in the attack landscape necessitates a deeper understanding of the current capabilities of malicious actors for utilizing AI to conduct cyber attacks.
@@
This study aims to investigate the potential of using state-of-the-art language AI in the context of spear phishing. There is a need to assess how current models fare in generating malicious messages for distribution to unknowing victims, and whether they currently perform on par or better than human counterparts. We also assess whether certain factors in a spear phishing message, such as personalization, tone, and word choice, contribute significantly to its effectiveness.
@@
The implications of this research are twofold. First, awareness about the malicious capabilities of current technologies can equip cybersecurity professionals with knowledge and understanding on how to safeguard against the risks posed by AI-generated spear phishing attacks. Second, through a deeper understanding of themes and elements associated with more deceptive spear phishing attacks, cybersecurity education can evolve in tandem with emerging threats.
@@ 2 Review of Related Literature
The body of literature related to spear phishing is expanding, with numerous studies investigating the factors that lead individuals to fall for phishing attacks [6, 9, 1, 10], as well as the application of AI for detecting and generating such attacks [8, 11].
Phishing, like other forms of social engineering, relies on various principles known as weapons of influence to manipulate targets into performing actions that benefit the attacker [9, 12]. These principles include reciprocation, which appeals to the target’s desire to repay favors (e.g., by offering a gift card); liking, which leverages the users’ tendency to comply with requests from people they like or perceive similarities with; scarcity, which exploits the target’s sense of urgency in response to limited resources or availability; social proof, which utilizes peer pressure to encourage compliance; authority, which appeals to the user’s inclination to follow authoritative entities, such as government agencies or law enforcement; commitment, which pressures individuals into adhering to their perceived obligations; and perceptual contrast, which presents two distinct scenarios and prompts the user to choose the more appealing option. Malicious messages often employ these principles to exploit potential victims.
As communication methods evolve, so too does the medium of phishing. Initially, phishing primarily relied on email as its mode of communication. Recently, however, phishing campaigns have begun to adopt alternative platforms such as SMS [13] and social media [14]. There is comparatively less literature available on these more specific modes of phishing attacks.
Several studies have explored the potential of applying machine learning (ML) for generating phishing attacks. Seymour and Tully employed a neural network to create spear phishing messages using data from trained social media posts [11], while Khan et al., utilized the now-outdated GPT-2 model to generate phishing attacks [15]. These studies, which tested their ML phishing attacks against humans, achieved varying degrees of success. Advancements in AI, have significantly improved the capability of large language models to generate human-like text. Models such as GPT-4 have demonstrated an unprecedented ability to produce coherent, relevant, and persuasive content, which can potentially be leveraged for malicious purposes [8, 16]. However, literature specifically addressing spear phishing attacks generated by LLMs remains limited, as research on using LLMs for crafting phishing messages has only recently emerged.
Comparing the effectiveness of human-generated versus AI-generated phishing attacks is a relatively unexplored area of research, yet one of growing importance. While humans can craft highly targeted and convincing phishing messages, AI-generated messages can achieve similar success rates with great potential to scale [8, 11]. As a result, studies that aim to increase education about AI manipulation are emerging [17]
With the scarcity of literature on spear phishing generated by LLMs, research to reach a better understanding of these risks is essential. While many studies offer updated training methods to inform employees and end-users about phishing attacks [18], the this study aligns with these efforts while also addressing the emerging threat of LLM-generated spear phishing. By evaluating how effectively publicly available LLMs can generate convincing spear phishing attacks, we can emphasize the need for countermeasures for potential malicious AI use. This study aims to bridge the gap in the existing literature and provide insights that can inform the development of more robust cybersecurity protocols.
@@ 3 Research Questions
We posit four research questions that this study aims to address to form a comprehensive understanding of the capabilities of LLM generated spear phishing SMS messages compared to human capabilities. We hereafter refer to the research questions in this paper as RQ1, RQ2, RQ3 and RQ4:
@@ 4 Methodology
This study comprises multiple steps aimed at simulating spear phishing attacks on human subjects using a novel methodology we call TRAPD, which stands for Threshold Ranking Approach for Personalized Deception. TRAPD is designed to analyze the perceived effectiveness of deceptive messages (e.g., spear smishing messages) that have been personalized to a specific target in an ethical manner. For this study, spear phishing messages were created and targeted to each subject, who then assessed whether the message would have been successful or not in a real life setting. First, personal information was gathered from willing human subjects. The anonymized personal information was then passed to human authors who crafted spear phishing messages specific to each human target. The same personal information was also used to create prompts to feed to the GPT-4 model to generate spear phishing messages. Both the human-made and the AI-made messages were then shown to each participant who rank ordered them based on their persuasiveness and evaluated each spear phishing “attack” based on whether it would 1) trick them, and 2) whether the message was made by a human or AI. The below section describes the phases of the TRAPD methodology that were conducted.
@@ 4.0 The TRAPD Methodology
One of the contributions of this paper is the introduction of the TRAPD methodology for evaluating personalized deceptive messages. While the focus of this paper has been on spear phishing SMS messages, we believe the methodology that we performed could be used to evaluate other types of deceptive content tailored to individuals (e.g., personalized disinformation, other types of spear phishing messages). We feel that TRAPD can provide valuable and unique insights. The methodology is designed to provide an ethical way to create and evaluate personalized deceptive messages. It was also designed to facilitate comparison of different messages, such as those created by humans versus AI, or those created about different topics.
@@
    Recruiting targets willing to share personal information with potential attackers (humans or AI).
@@
    Creating personalized deceptive messages aimed at the targets.
@@
    Having targets rank order (sort) the messages from most compelling to least compelling and choose a threshold above which they would be deceived.
@@
    Having targets provide qualitative assessments of their rationale for sorting messages.
@@
    (Optionally) Having targets “label” messages with a variable of interest (e.g., whether they believe a message was created by AI or not) and then provide qualitative assessments explaining their labeling.
The following subsections explain how we implemented the TRAPD methodology in this particular study. We hope it will serve as a template for future studies that use the method to evaluate other types of personalized deceptive content.
4.1 Recruiting targets
The first phase of research comprised recruiting participants as potential targets for spear phishing attacks. Recruiting included convenience sampling from people whom the researchers know, and using printed flyers and online banner advertising that directed interested individuals to complete a Qualtrics survey to sign up, consent to participate, and share some personal information for the purposes of this study. Participants were asked to participate in an experiment and post experiment interviews to provide detailed feedback on the spear phishing messages that were crafted for them. Participants were not told at recruitment how the spear phishing messages would be generated (i.e., human vs AI).
Participants were asked to share their name and email address for study contacting purposes. In addition, gender, age group, and self-reported technical ability were collected for study analysis. Additional information specific to each participant was also asked in order to generate personal references in the targeted spear phishing messages. This information included the participant hobbies, city or state in which they live, their occupation and workplace, and something from their home that they recently posted on social media.
After the sign-up period, we had 41 candidate participants who were all sent invitations. Twenty-five responded with complete registration information and arrived on their scheduled interview day. The participant demographics are shown on Table I.
TABLE I: Target Demographics
Gender Targets
Male 13 (52%)
Female 12 (48%)
Age Group Targets
18-25 9 (36%)
26-40 10 (40%)
40+ 6 (24%)
(a) Threshold Rank Ordering
(b) Label as Human vs AI
Figure 1: The two phases for the target interview.
@@ 4.2 Creating personalized deceptive messages
Using the personal data gathered from each participant, spear phishing messages were generated from both the GPT-4 model and from human authors. The goal of each message was to convince the human target to click the link included in each message. The spear phishing messages that were generated were in the form of SMS. For each target, we formed three different prompts, based on the information gathered during recruitment. Each prompt focused on one of three topics: the target’s hobbies, workplace, or something they own that they recently posted on social media. Below is an example of the prompt for one of the targets (name altered):
@@
This phase of work was conducted in two sub-phases, one for the human authors and one for GPT-4. The sub-phases below were conducted in parallel.
@@ 4.2.1 Human Generation
The human authors were undergraduate students enrolled in a university cybersecurity program during the Winter Semester 2023. Students generated spear phishing messages based on the data provided to them about their target, though names were changed to maintain participant privacy. Messages were entered into an online survey along with some demographic data about the authors including name, gender, age group, and technical ability. Authors were also asked to assess their understanding and experience relative to spear phishing. Authors were then provided with four prompts containing information about their target and then were asked to generate a message that would attempt to phish that target based on that information. After each prompt, the author was asked to assess their confidence in the ability of the message to trick the target.
Ninety-nine students participated in the survey, with the intent of creating a balanced pool of message contributions across demographics. In total, 297 messages were gathered and screened for invalid entries that were not spear phishing messages. On messages that have no link, or have placeholders for links, we modified the message to include a generic tinyurl. In cases where there are more than two valid messages for each topic (work, hobbies, social) for each target, two were selected at random to proceed to the next phase. 246 human-authored messages were retained and prepared for the next phase.

### v2

# spacing=nonfrench

@@ Abstract
This paper explores the use of Large Language Models (LLMs) in spear phishing message generation and evaluates their performance compared to human-authored counterparts. Our pilot study examines the effectiveness of smishing (SMS phishing) messages created by GPT-4 and human authors, which have been personalized for willing targets. The targets assessed these messages in a modified ranked-order experiment using a novel methodology we call TRAPD (Threshold Ranking Approach for Personalized Deception). Experiments involved ranking each spear phishing message from most to least convincing, providing qualitative feedback, and guessing which messages were human- or AI-generated. Results show that LLM-generated messages are often perceived as more convincing than those authored by humans, particularly job-related messages. Targets also struggled to distinguish between human- and AI-generated messages. We analyze different criteria the targets used to assess the persuasiveness and source of messages. This study aims to highlight the urgent need for further research and improved countermeasures against personalized AI-enabled social engineering attacks.
@@ 1 Introduction
Cybersecurity is a dynamic battle between attackers and defenders continuously adapting to new and emerging technologies. Although robust security measures can mitigate cyber risk, the strength of a system often depends on its users .[1] Among the myriad tactics exploited by threat actors, phishing remains the most common way to infiltrate systems. Phishing employs social engineering to persuade victims to perform an action, such as clicking on a malicious link or email attachment, that causes malicious code to run or disclose sensitive information. Phishing attacks usually take the form of email messages, telephone calls (vishing) or SMS messages (smishing) .[2] According to the FBI Internet Crime Complaint Center, phishing complaints in 2023 are five times more common than all other types of attacks, accounting for more than \$300 million in reported losses between 2021 and 2023 .[3]
@@
Spear phishing is a targeted phishing method that uses personalized messages with specific information, such as a user’s name, job title, or home address, making them harder to distinguish from legitimate communications .[4] Targeted spear phishing poses a significant threat for the future as its prevalence and sophistication continue to grow. According to the Proofpoint "State of the Phish 2023" report, spear phishing accounted for 74% of phishing attacks in 2022, compared to only 8% for bulk phishing .[5] In instances where the pretext used against a victim matches their expectation, the attacker is likely to succeed .[6] For instance, users expecting a shipping update from a retailer are more likely to fall for phishing messages imitating such communications .[7] The effectiveness of spear phishing stems from its personalization, a characteristic that AI technologies increasingly excel at replicating, but with far less effort than required by humans.
Traditionally, crafting these targeted messages was less common. It required significant time and effort by cybercriminals to research and develop messages specifically tailored to their victims .[7] However, with the advancement of Artificial Intelligence (AI), emergent large language models (LLMs) can be used to carry out spear phishing campaigns with greater efficiency and human-like accuracy .[8] LLMs like GPT-4 generate text that mimics human language, making them potentially powerful in crafting deceptive messages with greater speed and less resources. The ability of AI to create these messages that are almost indistinguishable from human-written messages has heightened the need to assess how cybersecurity defenses can account for these new capabilities. While human-generated phishing may rely on individual creativity and context, AI-generated messages can leverage scalability and efficiency.
@@
This study investigates the use of state-of-the-art language AI in crafting spear phishing SMS messages. We assess how current models fare in generating these personalized, malicious messages for distribution to potential targets and whether they currently perform on par or better than human creators. We also determine how well spear phishing targets can identify AI-generated messages, since such knowledge may be used to help protect oneself. Finally, we report qualitative feedback from targets on which factors of SMS spear phishing messages, such as personalization, tone, and word choice, contribute significantly to their effectiveness or failure. We do so with an eye toward the role of personalization, since it has the potential to both increase and decrease effectiveness.
@@
The implications of this research are two-fold. First, awareness of the malicious capabilities of current technologies can equip cybersecurity professionals with knowledge and understanding on how to protect against the risks posed by AI-generated spear phishing attacks. Second, through a deeper understanding of themes and elements associated with more deceptive spear phishing attacks, cybersecurity education can evolve in tandem with emerging threats. Furthermore, distinguishing AI-generated messages from human-generated ones could shape future policy and training aimed at individuals to more reliably identify AI-generated content.
@@ 2 Review of Related Literature
The body of literature related to spear phishing is expanding, with numerous studies investigating the factors that lead individuals to fall for phishing attacks [6, 9, 1, 10], as well as the application of AI for detecting and generating such attacks .[8, 11] However, a notable gap exists in understanding how effectively AI-generated messages compare to human-authored ones, particularly in the context of spear phishing and SMS phishing (i.e., smishing), as discussed in this section.
2.1 Evolution of Phishing Techniques
As communication methods evolve, so do the techniques and methods used in phishing attacks. Early phishing attacks primarily relied on email due to its widespread use and ease of access to consumers and businesses .[1, 12] These emails generally used simple deception techniques, such as forged email headers, deceptive subject lines, and language that tries to copy legitimate and authoritative entities like banks and commercial platforms. Less successful phishing attacks were easily recognized by these identifying elements, as well as egregious spelling or grammatical errors. But over time, these attacks became more sophisticated, with attacks containing content that is almost indistinguishable from legitimate sources. Eventually, attackers began incorporating personalized elements to make these attacks even more convincing in what became known as spear phishing. These attacks use messages with some information about specific individuals extracted from social media, public records, or previous breaches, making them appear much more credible. The shift towards more targeted attacks enabled malicious actors to execute more successful phishing campaigns.[4, 5]
While email remains a dominant vector for phishing attacks, phishing campaigns have begun to adopt alternative platforms such as SMS text messages [13] and social media .[14] SMS phishing, also known as smishing, takes advantage of our growing reliance on text messaging via mobile devices. Smishing targets users with shorter, but equally deceptive text messages that often appear to come from trusted entities like delivery services, financial institutions or online platforms. Since many legitimate companies use SMS to automate user alerts, attackers may forge these messages to prompt users to urgently "resolve" a fabricated issue.
Social media platforms are also becoming fertile ground for phishing attacks. Attackers can create fake profiles of individuals or organizations to reach out to their targets or publish deceptive posts or ads .[14, 15] These can lead unsuspecting victims to open malicious links or give sensitive information. The open nature of SMS and social networks makes it difficult to combat phishing in these mediums. Despite the diversity of these attack vectors, existing studies rarely address the extent to which message origin—–AI or human—–contributes to their success or detection. This suggests a need to evaluate the current threat landscape.
2.2 AI-enabled Phishing
The integration of machine learning (ML) and neural networks into phishing attacks has opened new avenues for cybercriminals. With the development of large language models (LLMs), it has become easier to apply deep learning techniques to tackle tasks that demand high cognitive ability and creativity. In a study by Zhai et al., models like ChatGPT and GPT-4 were used to solve a variety of tasks, ranging from common-sense reasoning to more complex problem solving .[16] The AI consistently outperformed non-college students in these assessments, showcasing the potential for AI to execute creative tasks, such as social engineering.
Seymour and Tully trained a neural network on social media posts to create spear phishing messages to showcase the potential of AI to create targeted attacks.[11] Their approach reported a success rate of 30-66%, compared to the 5-14% reported from large-scale phishing campaigns .[17] Khan et al., utilized the GPT-2 model to generate phishing attacks, although models continue to improve dramatically over time .[18] Despite this, they used a game-theoretic approach to improve upon the model so as to gain a competitive edge by calculating the risks and rewards before implementing the attack. The study illustrated the potential to combine other disciplines to conduct more successful campaigns against potential targets.
As large language models (LLMs) have become more commonplace, continuous advancements in AI have significantly improved the capability of LLMs to generate human-like text. Models like GPT-4 (and GPT-4o), Gemini, and Llama have demonstrated an unprecedented ability to produce coherent, relevant, and persuasive content, which can pose significant cybersecurity challenges.[8, 19] The literature on AI-generated spear phishing is still emerging, with limited research specifically focusing on the cybersecurity risks associated with LLMs. There is a significant gap in understanding whether AI-generated messages are inherently more or less effective than human-generated ones and how defenders can distinguish between the two. This comparison has implications for detection technologies and end-user education.
AI-driven tools, like deepfakes [20] or automated reconnaissance [21], can significantly amplify the capabilities of AI-generated phishing. Further investigation is needed to evaluate how AI-enabled attacks can evolve in combination with other technologies and disciplines and what countermeasures may be effective in mitigating these risks.
2.3 Comparison of Human vs AI in Phishing
Comparing the effectiveness of human-generated versus AI-generated phishing attacks is a relatively unexplored area of research, yet one of growing importance. The success of human-generated spear phishing attacks traditionally relied on the skill of the person in crafting convincing messages tailored to the target. Each human attacker differs in their ability to gather data from available sources, identify vulnerabilities in reaching their target, and their capacity to use social engineering to their advantage. Koivisto and Grassini showed that the best human ideas still outperforms the capabilities of AI in high-level, divergent creative tasks.[22] No matter the skill level, crafting these highly personalized spear phishing campaigns are time consuming and resource intensive. Thus, the primary limitation of human-generated spear phishing is the lack of scalability, as each message must be individually crafted to its target to maximize the chance of success.
On the other hand, AI-generated phishing presents a solution to that shortcoming—where LLMs can generate phishing messages at scale, with significantly less human input .[8, 11] State-of-the-art models can quickly produce content that can mimic human natural language, which means that attackers can potentially execute highly targeted phishing campaigns with great efficiency. Currently, AI-generated phishing attacks can lack the contextual nuances that a skilled human author may provide, but when combined with principles from other fields [18] or new data analysis techniques[11], AI-enabled attacks could eventually make up for this limitation. Studies have yet to establish whether AI-generated phishing attacks can systematically outperform human-generated ones, particularly in real-world settings. Addressing this research gap is crucial to understanding how both AI and human-generated attacks exploit human vulnerabilities. The probability that AI evolves to outperform human counterparts to create more enticing spear phishing messages is high. Even now, the current iteration of LLMs poses a growing threat that requires new detection and prevention methods .[23]
@@ 3 Research Questions
We posit four research questions to investigate the capabilities of LLM generated spear phishing SMS messages compared to human capabilities. We hereafter refer to the research questions in this paper as RQ1, RQ2, RQ3 and RQ4:
@@ 4 Methodology
This study comprises multiple steps using a novel methodology we call TRAPD, which stands for Threshold Ranking Approach for Personalized Deception. TRAPD is designed to analyze the perceived effectiveness of deceptive messages that have been personalized to a specific target (e.g., spear phishing messages) in an ethical manner. It uses quantitative methods to compare the perceived effectiveness of different types of messages (e.g., AI vs. human), as well as qualitative methods to understand what makes messages compelling (or not).
For this study, spear phishing SMS messages were created and evaluated by 24 human targets. Informed consent was received by the participants (i.e., targets), and the study was reviewed and approved by [anonymized IRB]. The following section describes the phases of the TRAPD Methodology, with details on how it was implemented in the current study to answer our research questions.
@@ 4.0 The TRAPD Methodology
One of the contributions of this paper is the introduction of the TRAPD methodology for evaluating personalized deceptive messages in an ethical manner. While the focus of this paper is on spear phishing SMS messages, the methodology may be used to evaluate other types of deceptive content tailored to individuals (e.g., personalized disinformation, other types of phishing messages). TRAPD is designed to facilitate comparison of different messages, such as those created by humans versus AI, or those created about different topics. Finally, it is designed to collect both quantitative and qualitative data, helping us to understand the "why" behind the quantitative findings.
@@
    Recruit targets who willingly share personal information with potential attackers.
@@
    Generate personalized deceptive messages aimed at the targets (e.g., using humans or AI).
@@
    Have targets rank order (sort) the messages from most compelling to least compelling and choose a threshold above which they would be deceived.
@@
    Have targets provide qualitative assessments of their rationale for placing messages where they did.
@@
    (Optionally) Having targets “label" messages with a variable of interest (e.g., whether they believe a message was created by AI or not) and then provide qualitative explanations for their labeling choices.
Figure 1 illustrates the flow of the current project, as described by these TRAPD methodology steps. The following subsections explain how we implemented the TRAPD methodology in this particular study. We hope it will serve as a template for future studies that use the method to evaluate other types of personalized deceptive content.
Figure 1: Overview of the project within the context of the TRAPD Methodology steps. Recruitment, Generation and Assessment is numbered according to each step in the methodology.
4.1 Recruiting targets who share personal information
The first phase of research comprised recruiting participants as potential targets for spear phishing attacks and collecting personal information from them as part of a survey. Recruiting included convenience sampling from people whom the researchers know, placing printed fliers in the university library, posting online fliers on personal social media accounts, and sending online fliers to email distribution lists of university departments and neighborhoods. We sought to recruit targets with a variety of jobs and different ages, rather than simply having students participate. Fliers directed interested individuals to complete a Qualtrics survey to sign up, consent to participate, and share some personal information for the purposes of this study. Participants were also told they would be invited to participate in an in-person interview where they would evaluate personalized messages created for them based on the information they provided. Participants were not told at recruitment time how the personalized messages would be generated (i.e., human vs AI).
A total of 41 people (i.e., targets) completely filled out the survey, which asked for their name, email address (to schedule the follow-up interview), sex, age group, and self-reported technical ability. Targets also provided details about themselves that would be used to generate the spear phishing messages including: participant hobbies, city and state in which they live, their occupation and workplace, and something from their home that they recently posted on social media.
@@ 4.2 Creating personalized deceptive messages
Using the personal data gathered from each participant, spear phishing messages were generated from both the GPT-4 model and from human authors. The goal of each message was to convince the human target to click the link included in each message. The spear phishing messages that were generated were in the form of SMS messages. Each prompt focused on one of three topics: the target’s hobbies, workplace, or something they own that they recently posted on social media. Below is an example of the prompt for one of the targets (name altered, as it was in the original prompts to maintain target’s anonymity):
@@
The same exact prompt was provided to human authors and GPT-4 as described in the following sections.
@@ 4.2.1 Human Generation
Human authors were recruited from undergraduate students enrolled in a university cybersecurity program or honors students taking a course on deception during Winter semester 2023. All students had learned about techniques used in phishing attacks and had practiced creating them as part of a class. They were given 15 minutes of time during classes to write up to 4 messages, each based on a different randomly selected user and prompt. Students received extra credit for their participation. Based on observations, students were highly invested in the task as evidenced by a lack of discussion or participation in other activities. A few stayed longer than the 15 minutes to finish their messages.
Authors typed their spear phishing messages into an online survey tool that allowed for emojis and had a text box that approximated the size/length of a standard SMS message. They were provided with 4 randomly selected prompts containing information about their target, then asked to generate an SMS spear phishing message that would use that information in a way that would get the user to click on a link. After each prompt, the author was asked to assess their confidence in the ability of the message to deceive the target. They also answered questions about their age group, sex, technical ability, and understanding and experience relative to spear phishing.
Ninety-nine student authors participated, with the intent of creating a balanced pool of message contributions across demographics. In total, 363 messages were gathered and screened for invalid entries that were not spear phishing messages or were obviously problematic (e.g., not completed). In cases where a user included only a placeholder for a link (e.g., "[link]") we replaced with a generic tinyurl link.
We expected that some human authors may not have created valid messages. For this reason, we collected at least 3 messages for each target/topic combination (e.g., target15/hobbie), even though we ultimately only showed targets 2 messages per topic. To give the human authors the best chance, all messages underwent an additional review by our team, which include two cybersecurity professors that teach courses that cover social engineering. We excluded any messages that were incomplete, did not include a prompt to click a link, or were of obviously low quality. In cases where there were more than two valid messages for each topic (work, hobbies, social) for each target, two were selected at random to proceed to the next phase. All in all, 246 human-authored messages were retained and prepared for the next phase. This process helped counter potential problems of having students be the authors of spear phishing messages.